OpenStudy (anonymous):

what is the percentage increase from 5000 to 15000?????

OpenStudy (anonymous):

300%

OpenStudy (anonymous):

explain please

OpenStudy (anonymous):

its 200% actually

OpenStudy (anonymous):

increase = 15000-5000 =10000

OpenStudy (anonymous):

original was 5000

OpenStudy (anonymous):

oops missed the increase bit

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[\frac{10000}{5000} \times 100=200%\]

OpenStudy (anonymous):

Thanks